# Clarifying identification and estimation of treatment effects in the Sequential Parallel Comparison Design

### v1

@@ 1 Introduction
Placebo response complicates statistical inference and clinical understanding regarding treatment efficacy in placebo-controlled clinical trials in psychiatry and other fields[fava2003, chen2011, rybin2018]. As a result, studies have less ability than expected to detect effective treatments due to estimating a smaller active treatment effect. Estimating placebo response is one way to correct for this bias in statistical inference of average active treatment effects. The sequential parallel comparison design (SPCD) was developed to address this gap by attempting to identify the placebo non-responders during the first stage and conducting the second stage only among those identified as non-responders[fava2003].
Typically participants are allocated to active and placebo arms at stage 1 using a 2:1 ratio with 2 participants assigned to placebo for each assigned to active treatment. A placebo response classification rule based on symptom improvement is then used to classify the placebo non-responders. Those classified non-responders are then randomized at stage 2 to active and placebo with roughly half assigned to each. In total this results in roughly a third of the participants receiving the active treatment at stage 1, another third receiving placebo and being categorized as placebo responders, and the final third being categorized as placebo non-responders and being re-randomized at stage 2. While alternative strategies to setting randomization allocation ratios are possible[tamura2011], the decision to choose a particular allocation is typically arbitrary. Additionally, the allocation scheme does not affect our results on the estimands and estimators in Section 3.
In the SPCD, participants are first randomized to active treatment or placebo at stage 1. After collecting outcomes for stage 1, participants in the placebo arm are classified as placebo non-responders or placebo responders. Placebo non-responders are re-randomized in stage 2 to receive placebo or active treatment. Figure 1 displays the trial design. Note that in this paper we use the term estimand, treatment effect, causal effect, and causal parameter interchangeably.
The intuition behind the design is that the placebo response in stage 1 dilutes the measured active treatment effect while the effect measured at stage 2 should only capture the active treatment effect without placebo response. The literature on statistical analysis for the SPCD proposes an overall estimate resulting from a weighted average of the stage 1 and stage 2 active treatment effect estimates[chen2011, rybin2018, tamura2011, chi2015, cui2019]. However, statistical inference under this design comes with significant complications[rybin2018, tamura2011, chi2015, rybin2015, liu2022]. In the rest of the manuscript, we use responders and non-responders to refer specifically to placebo response, unless otherwise noted. For brevity, we also use treatment effect or average treatment effect to refer specifically to the active treatment effect or average active treatment effect respectively.
@@
Prior work[chi2015] discussed six issues found with the conventional SPCD estimators. These issues include (1) placebo response biasing estimates, (2) assuming the active treatment effect is constant over time, (3) uncertainty in how to set the weight for the overall active treatment effect estimate averaging the stage 1 and stage 2 estimates, (4) unclear interpretation of the overall estimate, (5) inconsistency between the directions of the stage 1 and overall estimate, and (6) the common assumption that stage 1 and stage 2 are independent.
In this manuscript, we apply the counterfactual and causal inference framework[rubin1974] to illustrate the incoherence of the conventional weighted SPCD estimator, and to highlight an additional complication: placebo responder status misclassification. Others have discussed misclassification, primarily to motivate better predictive models[rybin2018, rybin2015, liu2022]. Here we show how crucial accurate placebo response classification is to accurately estimating the active treatment effects in the SPCD.
In Section 3, we develop quantities of interest, discuss identification assumptions required for estimation, and derive expectations for the conventional estimators in the SPCD using the causal inference framework. Using these derivations, we argue that the SPCD-specific estimands are not clearly defined and are either not targeted by the conventional estimators due to placebo response misclassification, or are targeted under unrealistic scenarios. Some of the unrealistic scenarios defeat the purpose of using an SPCD design (we provide more details on Sections 3.3.2, 3.3.3, and 5). The stage 1 estimand and estimator do target the average treatment effect for all participants, but this is achievable with a standard parallel randomized design. We demonstrate the importance of accurate placebo response classification using a simulation study in Section 4. Finally, in Section 5, we summarize our results, connect them to previously published work, and provide some direction on how to approach analysis for SPCD trials.
@@ 2.1 SPCD Set Up
We consider the binary latent placebo responder framework proposed by Rybin et al.[rybin2018] wherein the population is a mixture of placebo responders and non-responders. A variation of the latent responder status is also used more generally as a counterfactual to categorize both drug response and placebo response[liu2022]. Both latent responder approaches[rybin2018, liu2022] use an expectation maximization (EM) algorithm[dempster1977] to account for the unknown responder status when fitting their respective models.
@@ 2.1.1 Notation
For each of $i\in\{1,\ldots,n\}$ study participants, let $L_{i}$ denote a binary latent variable indicating whether a participant would respond to placebo ($L_{i}=1$) or not ($L_{i}=0$). Let $Y_{i,0}$, $Y_{i,1}$, and $Y_{i,2}$ denote the participant’s baseline (e.g., pre-treatment symptom severity), stage 1 and stage 2 outcomes, respectively. Let $A_{i,1}$ denote the randomized arm at Stage 1, i.e, placebo ($A_{i,1}=0$) or active treatment ($A_{i,1}=1$) and $A_{i,2}$ the assigned or randomized arm at stage 2, respectively. Finally, let $R_{i}$ represent the placebo-responder indicator, i.e., participants are classified as responders ($R_{i}=1$) or non-responders ($R_{i}=0$). The observed data are $D=(Z_{1},\ldots,Z_{n})$, where $Z_{i}$ represents the data for the $i^{th}$ participant, i.e., $Z_{i}=(Y_{i,0},Y_{i,1},Y_{i,2},R_{i},A_{i,1},A_{i,2})$ following some multivariate probability distribution $\P$. The latent exogenous variables $U_{Y_{0}}$, $U_{A_{1}}$, $U_{Y_{1}}$, and $U_{Y_{2}}$ are unmeasured and contain background information or influences on the observed variables[pearl2009causality].
@@ 2.1.2 A structural causal model and associated DAG
To clarify the SPCD set up, we use the structural causal model (SCM) and directed acyclic graph (DAG)[pearl2009causality] described in Figure 2. From the next paragraph, we omit the participant index $i$ for better readability.
@@ 3 SPCD Estimands and Estimators
We discuss potential estimands targeted in a SPCD in Section 3.1 as functions of counterfactual variables[pearl2009causality]. Then, we introduce assumptions for their identification in Section 3.2. Finally, we discuss conventional SPCD estimators in Section 3.3.
@@ 3.1 SPCD Estimands
Most previous literature[chen2011, rybin2018, tamura2011, chi2015, cui2019, rybin2015] defines the average treatment effect for all participants $\Delta_{\mathrm{all}}$ as the effect between treated participants and a mixture of placebo responders and non-responders. The treatment effects $\Delta_{NR}$ for the placebo non-responders, and $\Delta_{PR}$ for placebo responders are defined less often in the literature[rybin2018, chi2015]. We consider $\Delta_{PR}$ to be equivalent to $\Delta_{NR}-\Delta_{\mathrm{placebo}}.$ To our knowledge, previous work has not defined a quantity equivalent to $\Delta_{\mathrm{placebo}},$ the size of the placebo response for placebo responders relative to placebo non-responders.
@@
Average Placebo Response Effect $\Delta_{\mathrm{placebo}}=E(Y_{1}^{a_{1}=0,l=1})-E(Y_{1}^{a_{1}=0,l=0})$ Stage 1 Placebo-treated Participants; $A_{t}=0$
@@ 3.2 Identification Assumptions
Assumption 1 states that each unit’s observed outcome corresponds to their latent placebo response status and their counterfactual outcome under the treatment actually received and that treatment is consistently applied or given across individuals. In other words, there are not different versions of the placebo or active treatment due to variation in providers, dosage, adherence, etc. It also assumes no interference between units. Together these comprise the stable unit treatment values assumption[rubin1980, robins2000, hernan2020]. This is encoded in the structural causal model (Figure 2).
@@
If $\mu_{t}^{a=1,l=1}\neq\mu_{t}^{a=1,l=0},$ then we are unable to identify the ATE for non-responders ($L=0$) because we do not classify active treatment participants as responders/non-responders in order to be able to estimate $\mu_{t}^{a=1,l=0}.$ While this assumption could be violated in practice, we make it here for brevity in the exposition of the issues with the estimators. Alternative strategies that do not require this assumption have been proposed but are beyond the scope of this paper[liu2022].
@@
Assumption 6 imposes a parametric mixture structure, as in Equation 6 which enables model-based identification: if the mapping $(p_{L},\xi_{0},\xi_{1})\;\longmapsto\;f_{Y_{1}\mid A_{1}=0}(\cdot)$ is one-to-one within the chosen family $\mathcal{F}$, then $(p_{L},\xi_{0},\xi_{1})$ and thus $(\mu_{1}^{0,0},\mu_{1}^{0,1})$ are identified under that model. Note this identifies component means, but does not imply we can correctly classify individual $L$’s. While other identification strategies for causal inference with latent variables have been proposed (see, e.g., [zhu2024causal, zhang2021bounding]), in this manuscript we adopt a model-based approach. As our goal is to clarify causal effects in SPCD trials, the model-based framework suffices for that purpose, as shown in the following sections and in the Supplementary Materials.
@@
Based on these assumptions, we provide identification of Table 1 estimands in Appendix  5 and that will be used to clarify what estimands are targeted by the conventional SPCD estimators.
@@ 3.3 The Conventional SPCD Estimators
The commonly used weighted estimator[chen2011, rybin2018, tamura2011, chi2015, cui2019, liu2022] combines stage 1, $\hat{\theta}_{1}$, and stage 2, $\hat{\theta}_{2}$, to get $\hat{\theta}_{w}$ using an pre-determined scalar weight $w$. See Appendix 5 for full derivations of the expectations.
@@ 3.3.3 Overall Weighted Estimator
A similar latent responder status framework could be used for active treatment response such that participants may or may not respond to the active treatment. However, active treatment non-responders are then effectively receiving a placebo and so are part of the placebo responder status framework as well. This is a key part of the latent responder status framework[liu2022].
@@ 3.4 The (Lack of) Attention Given to Classification
A limitation of previous work on SPCD is the absence of substantial discussion of placebo response classification, with limited exceptions[rybin2018, rybin2015, liu2022]. Typically, classification is done by setting a cut-off or threshold $c$ as in Equation 15,
@@
so that participants with large improvements over baseline or large outcomes in the placebo arm at Stage 1 are labelled as placebo responders. As previously noted[rybin2018, rybin2015, liu2022], this approach is subjective due to the arbitrary choice of $c$, error-prone, and often not supported by any relevant clinical data. However, given that the placebo response status is unobservable, this is a relatively intractable classification problem.
One attempt[rybin2015] at addressing the misclassification issue proposed using a re-randomization design that re-assigns treatment to all Stage 1 placebo participants rather than just placebo non-responders and then weights their response in Stage 2 by a measure of their responder status. Another approach[rybin2018] went further in the discussion of misclassification and proposed using a normal mixture model to classify the responders and non-responders.
This mixture model approach was expanded upon by Liu et al. (2022)[liu2022] and the misclassification issue was given more extensive discussion. They note perfect classification as being a crucial assumption in the derivation of their strata-level effects, estimators, and inference. Liu et al.[liu2022] also criticize the use of thresholds as subjective and error-prone which can potentially bias the inference. Instead they also propose fitting a mixture model with an EM algorithm and using the fitted probabilities of belonging to each strata given the data to then sample the “observed” responder statuses $S_{i}^{a_{1}}$ and $S_{i}^{a_{2}}$ for each individual $i.$ However, the sampled responder statuses are still treated as known. They discuss how an approach that treats these statuses as stochastic may be an avenue for further development. The stochastic approach is beyond the scope of this paper. Liu et al.’s simulation studies[liu2022] demonstrate that the causal estimators derived through their approach are biased when assuming the sampled placebo response are fixed. This is likely due to the misclassification which they note in their discussion.

### v2

# Clarifying the role of placebo response classification in the analysis of the Sequential Parallel Comparison Design

@@ 1 Introduction
Placebo response complicates statistical inference and clinical understanding of treatment efficacy in placebo-controlled clinical trials in psychiatry and other fields[6, 1, 15]. As a result, studies have less ability than expected to detect effective treatments due to estimating a smaller active treatment effect. Estimating placebo response is one way to correct for this bias in statistical inference of average active treatment effects. The sequential parallel comparison design (SPCD) was developed to address this gap by attempting to identify the placebo non-responders during the first stage of the design, and conducting the second stage only among those identified as non-responders[6].
Typically participants are allocated to active and placebo arms at stage 1 using a 2:1 ratio with 2 participants assigned to placebo for each assigned to active treatment. A placebo response classification rule based on symptom improvement is then used to classify the placebo non-responders. Those classified as non-responders after stage 1 are then randomized in stage 2 in equal ratio to active treatment or placebo. In total this often results in roughly a third of the participants receiving the active treatment at stage 1, a third receiving placebo and being categorized as placebo responders, and the final third being categorized as placebo non-responders and being re-randomized at stage 2. While alternative strategies to setting randomization allocation ratios are possible[18], the decision to choose a particular allocation is typically arbitrary. Additionally, the allocation scheme does not affect our results regarding the estimands and estimators in Section 3. Figure 1 displays the trial design. Note that in this paper we use the term estimand, treatment effect, causal effect, and causal parameter interchangeably.
The intuition behind the design is that the placebo response in stage 1 dilutes the measured active treatment effect while the effect measured at stage 2 should only capture the active treatment effect without placebo response. The literature on statistical analysis for the SPCD proposes an overall estimate resulting from a weighted average of the stage 1 and stage 2 active treatment effect estimates[1, 15, 18, 2, 3]. However, statistical inference under this design comes with significant complications[15, 18, 2, 14, 9]. In the rest of the manuscript, we use responders and non-responders to refer specifically to placebo response, unless otherwise noted. For brevity, we also use treatment effect or average treatment effect to refer specifically to the active treatment effect or average active treatment effect respectively.
@@
Prior work[2] discussed six issues found with the conventional SPCD estimators. These issues include (1) placebo response biasing estimates, (2) assuming the active treatment effect is constant over time, (3) uncertainty in how to set the weight for the overall active treatment effect estimate averaging the stage 1 and stage 2 estimates, (4) unclear interpretation of the overall estimate, (5) inconsistency between the directions of the stage 1 and the overall estimate, and (6) the common assumption that stage 1 and stage 2 are independent.
In this manuscript, we apply the counterfactual and causal inference framework[12] to illustrate the incoherence of the conventional weighted SPCD estimator, and to highlight an additional complication: placebo responder status misclassification. Others have discussed misclassification, primarily to motivate better predictive models[15, 14, 9]. Our approach to developing estimands and estimators using causal inference tools, discussed in Section 3, is similar to, but distinct and independently developed from those proposed in recent work by Liu et al.[9] and by Song et al.[16].
In particular, the recent work by Song et al.[16] proposes a formulation in which placebo responder status is an unobserved baseline trait, and the usual SPCD responder/non-responder classification based on the stage 1 outcome is treated as an imperfect surrogate. Their causal estimands for stage 1 are identified under assumptions that are coherent in this latent-baseline framework in which baseline covariates fully explain the association between responder status and outcomes, and that the probability of being a responder given baseline covariates can be correctly modeled. However, these assumptions do not hold for the SPCD as actually implemented. In practice, responder status is defined only after observing the stage 1 outcome. In contrast, we adopt a structural causal model that respects the operational SPCD classification rule, and we show that under this model these and related treatment effects are not nonparametrically identified without additional mixture assumptions, and that the conventional weighted SPCD estimator does not target a well-defined causal estimand. We provide additional discussion of this point in our analysis of the ADAPT-A trial in Section 5, where we argue that the latent-baseline formulation requires implausibly strong assumptions and reflects a clear mismatch with the SPCD design as actually implemented. In addition, we show how crucial accurate placebo response classification is to accurately estimating the active treatment effects in the SPCD.
In Section 3, we develop quantities of interest, discuss identification assumptions required for estimation, and derive expectations for the conventional estimators in the SPCD using the causal inference framework. Using these derivations, we argue that the SPCD-specific estimands are not clearly defined and are either not targeted by the conventional estimators due to placebo response misclassification, or are targeted under unrealistic scenarios. Some of the unrealistic scenarios defeat the purpose of using an SPCD design (we provide more details on Sections 3.3.2, 3.3.3, and 6). The stage 1 estimand and estimator do target the average treatment effect for all participants, however this is achievable with a standard parallel randomized design. We demonstrate the importance of accurate placebo response classification using a simulation study in Section 4 and with an illustration using a depression trial example in Section 5. Finally, in Section 6, we summarize our results, connect them to previously published work, and provide some direction on how to approach analysis for SPCD trials.
@@ 2.1 SPCD Set Up
We consider the binary latent placebo responder framework proposed by Rybin et al.[15] wherein the population is a mixture of placebo responders and non-responders. A variation of the latent responder status is also used more generally as a counterfactual to categorize both drug response and placebo response[9]. Both latent responder approaches[15, 9] use an expectation maximization (EM) algorithm[4] to account for the unknown responder status when fitting their respective models.
@@ 2.1.1 Notation
For each of $i\in\{1,\ldots,n\}$ study participants, let $L_{i}$ denote a binary latent variable indicating whether a participant would respond to placebo ($L_{i}=1$) or not ($L_{i}=0$). Let $Y_{i,0}$, $Y_{i,1}$, and $Y_{i,2}$ denote the participant’s baseline (e.g., pre-treatment symptom severity), stage 1 and stage 2 outcomes, respectively. Let $A_{i,1}$ denote the randomized arm at Stage 1, i.e, placebo ($A_{i,1}=0$) or active treatment ($A_{i,1}=1$) and $A_{i,2}$ the assigned or randomized arm at stage 2, respectively. Finally, let $R_{i}$ represent the placebo-responder indicator, i.e., participants are classified as responders ($R_{i}=1$) or non-responders ($R_{i}=0$). The observed data are $D=(Z_{1},\ldots,Z_{n})$, where $Z_{i}$ represents the data for the $i^{th}$ participant, i.e., $Z_{i}=(Y_{i,0},Y_{i,1},Y_{i,2},R_{i},A_{i,1},A_{i,2})$ following some multivariate probability distribution. The latent exogenous variables $U_{Y_{0}}$, $U_{A_{1}}$, $U_{Y_{1}}$, and $U_{Y_{2}}$ are unmeasured and contain background information or influences on the observed variables[10].
@@ 2.1.2 A structural causal model and associated DAG
To clarify the SPCD set up, we use the structural causal model (SCM) and directed acyclic graph (DAG)[10] described in Figure 2. From the next paragraph, we omit the participant index $i$ for better readability.
@@ 3 SPCD Estimands and Estimators
We discuss potential estimands targeted in a SPCD in Section 3.1 as functions of counterfactual variables[10]. Then, we introduce assumptions for their identification in Section 3.2. Finally, we discuss conventional SPCD estimators in Section 3.3.
@@ 3.1 SPCD Estimands
Most previous literature[1, 15, 18, 2, 3, 14] defines the average treatment effect for all participants $\Delta_{\mathrm{all}}$ as the effect between treated participants and a mixture of placebo responders and non-responders. The treatment effects $\Delta_{NR}$ for the placebo non-responders, and $\Delta_{PR}$ for placebo responders are defined less often in the literature[15, 2]. We consider $\Delta_{PR}$ to be equivalent to $\Delta_{NR}-\Delta_{\mathrm{placebo}}.$ To our knowledge, previous work has not defined a quantity equivalent to $\Delta_{\mathrm{placebo}},$ the size of the placebo response for placebo responders relative to placebo non-responders.
@@
Average Placebo Response Effect $\Delta_{\mathrm{placebo}}=E(Y_{1}^{a_{1}=0,l=1})-E(Y_{1}^{a_{1}=0,l=0})$ Stage 1 Placebo-treated; $A_{t}=0$
@@ 3.2 Identification Assumptions
Assumption 1 states that each unit’s observed outcome corresponds to their latent placebo response status and their counterfactual outcome under the treatment actually received and that treatment is consistently applied or given across individuals. In other words, there are not different versions of the placebo or active treatment due to variation in providers, dosage, adherence, etc. It also assumes no interference between units. Together these comprise the stable unit treatment values assumption[13, 11, 8]. This is encoded in the structural causal model (Figure 2).
@@
If $\mu_{t}^{a=1,l=1}\neq\mu_{t}^{a=1,l=0},$ then we are unable to identify the ATE for non-responders ($L=0$) because we do not classify active treatment participants as responders/non-responders in order to be able to estimate $\mu_{t}^{a=1,l=0}.$ While this assumption could be violated in practice, we make it here for brevity in the exposition of the issues with the estimators. Alternative strategies that do not require this assumption have been proposed but are beyond the scope of this paper[9].
@@
Assumption 6 imposes a parametric mixture structure, as in Equation 6 which enables model-based identification: if the mapping $(p_{L},\xi_{0},\xi_{1})\;\longmapsto\;f_{Y_{1}\mid A_{1}=0}(\cdot)$ is one-to-one within the chosen family $\mathcal{F}$, then $(p_{L},\xi_{0},\xi_{1})$ and thus $(\mu_{1}^{0,0},\mu_{1}^{0,1})$ are identified under that model. Note this identifies component means, but does not imply we can correctly classify individual $L$’s. While other identification strategies for causal inference with latent variables have been proposed (see, e.g., [22, 21]), in this manuscript we adopt a model-based approach. As our goal is to clarify causal effects in SPCD trials, the model-based framework suffices for that purpose, as shown in the following sections and in the Supplementary Materials.
@@
Based on these assumptions, we provide identification of Table 1 estimands in Appendix  Estimand Derivations and that will be used to clarify what estimands are targeted by the conventional SPCD estimators.
@@ 3.3 The Conventional SPCD Estimators
The commonly used weighted estimator[1, 15, 18, 2, 3, 9] combines stage 1, $\hat{\theta}_{1}$, and stage 2, $\hat{\theta}_{2}$, to get $\hat{\theta}_{w}$ using an pre-determined scalar weight $w$. See Appendix Estimator Derivations for full derivations of the expectations.
@@ 3.3.3 Overall Weighted Estimator
A similar latent responder status framework could be used for active treatment response such that participants may or may not respond to the active treatment. However, active treatment non-responders are then effectively receiving a placebo and so are part of the placebo responder status framework as well. This is a key part of the latent responder status framework[9].
@@ 3.4 The (Lack of) Attention Given to Classification
A limitation of previous work on SPCD is the absence of substantial discussion of placebo response classification, with some recent exceptions[15, 14, 9, 16]. Typically, classification is done by setting a cut-off or threshold $c$ as in Equation 15,
@@
so that participants with large improvements over baseline or large outcomes in the placebo arm at Stage 1 are labeled as placebo responders. As previously noted[15, 14, 9], this approach is subjective due to the arbitrary choice of $c$, error-prone, and often not supported by any relevant clinical data. However, given that the placebo response status is unobservable, this is a relatively intractable classification problem[16].
One attempt[14] at addressing the misclassification issue proposed using a re-randomization design that re-assigns treatment to all Stage 1 placebo participants rather than just placebo non-responders and then weights their response in Stage 2 by a measure of their responder status. Another approach[15] went further in the discussion of misclassification and proposed using a normal mixture model to classify the responders and non-responders.
This mixture model approach was expanded upon by Liu et al. (2022)[9] and the misclassification issue was given more extensive discussion. They note perfect classification as being a crucial assumption in the derivation of their strata-level effects, estimators, and inference. Liu et al.[9] also criticize the use of thresholds as subjective and error-prone which can potentially bias the inference. Instead they also propose fitting a mixture model with an EM algorithm and using the fitted probabilities of belonging to each strata given the data to then sample the “observed” responder statuses $S_{i}^{a_{1}}$ and $S_{i}^{a_{2}}$ for each individual $i.$ However, the sampled responder statuses are still treated as known. They discuss how an approach that treats these statuses as stochastic may be an avenue for further development. The stochastic approach is beyond the scope of this paper. Liu et al.’s simulation studies[9] demonstrate that the causal estimators derived through their approach are biased when assuming the sampled placebo response are fixed. This is likely due to the misclassification which they note in their discussion.
@@ 4 Simulation Study
5 Example: ADAPT-A Benchmark
As noted by Song et al.[16], the ADAPT-A trial[7] has served as a benchmark for methodological developments in the analysis of the SPCD[17, 1, 5, 15, 16]. Given its status as a benchmark data set, we defer extensive description and discussion of the data to its original publication[7] and other statistical methodology work that developed previous estimators[1, 5, 15, 16]. In this section, we present our estimates of the conventional SPCD estimands for the ADAPT-A trial using our unadjusted estimators proposed in Section 3. Alongside our estimates, we also present estimates using previously published methods including seemingly unrelated regression (SUR)[17], ordinary least squares (OLS)[1], linear mixed models (MIX)[5], imputation within an EM algorithm (EMIMP)[15], and the unadjusted (UNADJ (RC)) and OLS (OLS (RC)) methods following the cross-validated placebo response reclassification proposed by Song et al.[16] (where “RC" stands for reclassification). Note that we do not include the estimators from Liu et al.[9] nor the G-computation, IPW, or doubly robust estimators from Song et al.[16] as these estimators do not target the same quantities as the previously proposed conventional estimators.
Interpreting the ADAPT-A trial within the latent-baseline framework of Song et al.[16] would require positing an unobserved baseline placebo-response label and assuming that it is fully captured by baseline covariates, with the stage 1 responder classification acting as a simple surrogate. In ADAPT-A, however, the only operational definition of “responder” is a post-treatment function of the stage 1 outcome, and no baseline variables plausibly exhaust the drivers of placebo response. Under a structural causal model that respects this actual SPCD mechanism, the baseline responder label and its propensity are not identified from the data without strong, unverifiable modeling assumptions; thus the latent-baseline framework is not attainable for this trial. In other words, the assumptions required by this latent-baseline formulation are implausibly strong for ADAPT-A and reflect a clear mismatch with the SPCD design as actually implemented.
Table 2: Estimates for the ADAPT-A trial benchmark data set from [7, 16]. Our proposed unadjusted method is labeled “UNADJ". The other methods and estimates were previously published: “SUR" is seemingly unrelated regression from Tamura et al.[17], “OLS" is ordinary least squares regression from Chen et al.[1], “MIX" is a linear mixed model from Doros et al.[5], “EMIMP" is the EM algorithm-based estimation from Rybin et al.[15], and “UNADJ (RC)" and “OLS (RC)" are the unadjusted and ordinary least squares estimates using the placebo response reclassification step proposed by Song et al.[16].
Method $\hat{\Delta}_{all}$ $SE(\hat{\Delta}_{all})$ $\hat{\Delta}_{NR}$ $SE(\hat{\Delta}_{NR})$ $\hat{\Delta}_{w=0.75}$ $SE(\hat{\Delta}_{w=0.75})$ P-value
SUR -0.526 0.955 -2.228 1.125 -0.952 0.769 0.216
OLS -0.532 1.238 -2.173 1.126 -0.942 0.970 0.331
OLS (RC) -0.532 1.164 -1.914 1.053 -0.878 0.912 0.336
MIX -0.527 1.233 -2.382 1.132 -0.991 0.967 0.306
EMIMP -0.447 1.146 -2.353 1.493 -0.924 0.937 0.324
UNADJ -0.447 1.195 -2.197 1.112 -0.885 0.938 0.346
UNADJ (RC) -0.447 1.168 -2.008 1.077 -0.837 0.917 0.361
Let $\hat{\theta}_{t}^{M}$ be the point estimate at stage $t$ with method $M$ from the set of methods $\{UNADJ,OLS,SUR,MIX,EMIMP\}$, and let $SE(\hat{\theta}_{t}^{M})$ be the standard error for that point estimate. For the methods SUR, OLS, MIX, and EMIMP, the standard errors were calculated using closed form estimators. For the UNADJ, UNADJ (RC), and OLS methods, the standard errors were approximated using a bootstrap with 1,000 replicates.
The estimates from each of these methods are displayed in Table 2. Note that the stage 1 estimates $\hat{\theta}_{1}^{M}$ from the unadjusted or OLS estimation methods are unaffected by the placebo response reclassification proposed by Song et al., i.e. $\hat{\theta}_{1}^{UNADJ}=\hat{\theta}_{1}^{UNADJ~(RC)}=-0.447$ and $\hat{\theta}_{1}^{OLS}=\hat{\theta}_{1}^{OLS~(RC)}=-0.532$. Additionally, the unadjusted stage 1 point estimate is also equivalent to the stage 1 point estimate with the EM algorithm, $\hat{\theta}_{1}^{UNADJ}=\hat{\theta}_{1}^{EMIMP}=-0.447$. However, all of the standard errors for these point estimates vary due to being calculated analytically or approximated via bootstrapping. The regression based methods (OLS, OLS, SUR, and MIX) all produced stage 1 point estimates of approximately -0.530.
The range of the stage 2 estimates ($\hat{\theta}_{2}^{MIX}=-2.382$ to $\hat{\theta}_{2}^{OLS~(RC)}=-1.917$) is greater than the range of the stage 1 estimates ($\hat{\theta}_{1}^{OLS}=-0.532$ to $\hat{\theta}_{1}^{UNADJ}=-0.447$). The additional variation we see in stage 2 point estimates could be due to reducing the sample size at stage 2, or due to differences in how placebo response classification is addressed. In particular, the differences in placebo response classification due to the reclassification proposed by Song et al. results in the unadjusted and OLS estimates to be attenuated.
The large standard errors at both stages result in large standard errors for the SPCD estimator $\hat{\theta}_{w}^{M}$ for all of the methods as well. This results in no estimates being statistically significant at the $\alpha=0.05$ level for the hypothesis test of $H_{0}:\Delta_{w=0.75}=0$ vs $H_{1}:\Delta_{w=0.75}\neq 0$. The p-values were calculated using a normal approximation $p=2\left(1-\Phi\left(|\hat{\Delta}_{w=0.75}|~/~SE(\hat{\Delta}_{w=0.75}\right)\right)$.